determine if the system has no solutions, one solutions, or many solutions.
2x+y=-8
-8x-4y=32

victoria_is_hawt@yahoo.com

Answer by user_dude2008(1862)

2x+y=-8
-8x-4y=32


8x+4y=-32
-8x-4y=32


0x+0y=0

0=0 <---- True

Answer: infinite solutions